OpenStudy (anonymous):

Determine the geometric mean of 1, 3 , 5, 7, 9. Round your answer to the hundredth place. a. 2.94 c. 4.94 b. 1.94 d. 3.94

OpenStudy (anonymous):

alright

hartnn (hartnn):

for geometric mean, we need to take n'th root of the product of all terms n being number of terms \(here, \Large \sqrt[5]{1\times 3 \times 5\times 7\times 9}\)

OpenStudy (anonymous):

t im confused, i dont add then?

hartnn (hartnn):

nopes, for arithmetic mean, you add for geometric mean, you multiply :)

hartnn (hartnn):

you tried to help, thats already a good thing :)

OpenStudy (anonymous):

hahah okay then :) thanks

OpenStudy (anonymous):

i got 945 when i multipled them

hartnn (hartnn):

take the 5th root now :)

OpenStudy (anonymous):

how do i do that? o.o

hartnn (hartnn):

use calculator or google you should get 3.936 ...rounded to 3.94 :)

OpenStudy (anonymous):

how do i put that on my caculator though? haha thanks !

hartnn (hartnn):

welcome ^_^